# Incommensurable matter-wave jets in quasi-1D geometry

### v3

@@ Abstract
Matter-wave jets occur when a BEC is subjected to a modulation of interaction [21]. The modulation leads to a the formation of density waves inside the BEC followed by the emission of jets [22]. The microscopic process behind jet emission is photon-stimulated two-atom collisions [21]. First-order jets form from pairs of atoms that each absorb half of the energy of the photon. Further collisions lead to higher-order jets having velocities that can either be rational multiples of first-order jet velocity (commensurate jets) or not (incommensurate jets).
@@
Here we experimentally and numerically demonstrate the formation of incommensurable “golden” jets and other higher-order matter-wave jets in a quasi-1D BEC under a single-frequency interaction modulation. The incommensurable golden jets are precursors of the supercontinuum formation, which remains to be experimentally confirmed.
@@ I.1 Formation of matter-wave jets
Figure 2: (a) The final density profile ($t=40$) and (b) the time evolution of the density profile of the simulation for $\omega=33$, $k_{\mathrm{dc}}=-0.56$, $k_{\mathrm{ac}}=27.1$, $t_{\mathrm{p}}=8\pi$ and the fraction of atoms in different jets in (b) linear and (c) logarithmic scale. The horizontal line at 0.006 represents the threshold used to construct the phase diagram.
Figure 3: (a) Phase diagram of jets from GPE simulations. Different colors present regions where different types of jets are above threshold. (b) Final density profiles in momentum space $|\psi(q)|^{2}$ ($t=30$) for $\omega=33$ [along the dotted line in (a)]. (c) and (e) Experimental phase diagram of jets, measured at $t_{p}=$ 30 ms and 10 ms. The grey lines indicate where the number of atoms falls by 25% and 50% from the initial number of atoms in the condensate. The dashed black line in (c) is a fit to the boundary between the ”no jets” and ”J1” region, proportional to $\omega^{1/2}$. The errorbars correspond to the variance of the atom number and how precisely we can determine the amplitude corresponding to the boundary between regions. (d) and (f) Averages of 50 absorption images, as examples of different points in the phase diagram (marked with black crosses1 in the diagrams above). From bottom to top the images represent regions with: no jets, only J1, J1 and J2, and J1, J2 and G2 (top two images). The images were taken for evolution time $t_{e}=$ 42.7 ms. $k_{\mathrm{ac}}=Na_{\mathrm{ac}}/a_{r}$, where $N=4000\pm 200$ is the number of atoms in the BEC, before we start modulation. $a_{\mathrm{ac}}$ is calculated from the amplitude of the pulse by a calibration determined using the Feshbach resonance (see Methods).
@@ I.2 Numerical simulation of Gross-Pitaevskii equation
Based on calculated atom fractions, we construct a phase diagram of visibility of different sets of jets, shown in Fig. 3(a). A new region is declared to exist when the fraction for the next jet exceeds 0.006. We choose this value so that it is much larger than the noise in the simulation, but smaller than the atom fraction of the jets at high modulation amplitudes. As the amplitude $k_{\mathrm{ac}}$ increases, additional jets appear, represented by new regions in the phase diagram. As discussed above the transitions are actually continuous and not discrete, except from the “no jets” to the “J1” region. G3 jets occur in parts of the region marked with ”+G2” in Fig. 3(a).
@@ I.3 Experimental phase diagram of jet formation
In contrast to simulation, in the experiments, the number of atoms in jets varies from shot to shot; for example, see Fig. 1(a). This was previously observed in 2D [27], where the distribution of population in angular modes resembles the distribution of thermal radiation. More information on the statistics of the number of atoms in each type of jet is available in the Supplemental Material (Fig. S3). Because of this variance, gathering statistics is necessary to determine the presence of the phase at each point in the phase diagram. We choose the boundary to be at the point where at least one image out of ten shows the corresponding jets. Examples of averages of 50 absorption images of jets from different parts of the phase diagram are shown in Fig. 3(d) and (f) and in the Supplemental Material (Fig. S2). Note that the distance between the jets and the initial condensate varies for different modulation amplitudes. This is because jets are emitted sooner if the modulation amplitude is higher.
The range of frequencies in the experimental and the theoretical phase diagram are similar (the frequencies shown in Fig.  3(a) range from 0 to 7.2 kHz (from 0 to 80 in units of $\omega_{r}=2\pi\cdot 90\mathrm{Hz}$), whereas the experimental phase diagrams in Fig.  3(c) and (e) were measured for frequencies between 1 kHz and 7 kHz). On the other hand, the values of $k_{\mathrm{ac}}$ required for excitation of jets are approximately two times lower in the experiment than in simulation. The difference is most probably because the simulation describes a true 1D system, while the experiment, the BEC is a 3D object with jets confined into quasi-1D system by a dipole trap beam.
@@
As we increase the amplitude of modulation in the experiment, the total number of atoms starts decreasing and the incoherent background between the jets becomes stronger. This can be seen in Fig. 3, where we plot the thresholds at which the total number of atoms is 25% and 50% smaller than the number of atoms in the initial BEC. The decrease in the number of atoms is likely due to partial collapses of the BEC, caused by the strong attractive interactions the BEC is subjected to during the modulation leading to three-body losses. The incoherent background is formed by the excited atoms that are not part of the jets. Their velocities have random directions and that appears as a homogeneous background between J0 and J1, and to a lesser extent between J1 and J2. The momentum distribution of the incoherent background remains of the same width for increasing amplitude and does not increase beyond the position of J2 jets as the supercontinuum would. Atom losses and increased fraction of incoherent background, both not accounted for in the simulation, decrease the effective modulation amplitude and prevent the supercontinuum from forming.
@@ III.2 Numerical simulation
in the dimensionless form [30], where time $t$ is given in units $1/\omega_{r}$, $\omega$ in units $\omega_{r}$, and distances in units $a_{r}=\sqrt{\hbar/m\omega_{r}}$. $k$ is the interaction parameter $k(t)=Na(t)/a_{r}$. We solve the equation on an uniform 1D grid of 65536 points in the interval [-1200:1200]. The initial state of the BEC is determined by an imaginary-time evolution, while the calculation of the real-time dynamics is performed with a time-step of $\tau=$ 0.001. The code is written in Wolfram Language (Mathematica) and it can solve for the ground state through imaginary-time propagation of an initial approximation, and determine the time-dependence in real-time. Most of the computation time is spent in performing the (inverse) Fourier transform, which is performed highly efficiently in Mathematica, thus the performance of the code is comparable to that of compiled codes while it is significantly more flexible. Further details and examples can be found on: http://auger.ijs.si/gpe/.

### v4

@@ Abstract
Matter-wave jets occur when a BEC is subjected to a modulation of interaction [21]. The modulation leads to the formation of density waves inside the BEC followed by the emission of jets [22]. The microscopic process behind jet emission is photon-stimulated two-atom collisions [21, 23], where a quantum of the magnetic field modulation energy is transferred to the atom pair. First-order jets form from pairs of atoms that each absorb half of the energy of the photon. Further collisions lead to higher-order jets having velocities that can either be rational multiples of first-order jet velocity (commensurate jets) or not (incommensurate jets).
@@
Here, we experimentally and numerically demonstrate the formation of incommensurable “golden” jets and other higher-order matter-wave jets in a quasi-1D BEC under a single-frequency interaction modulation. The incommensurable golden jets are precursors of the supercontinuum formation, which remains to be experimentally confirmed.
@@ I.1 Formation of matter-wave jets
Figure 2: (a) The final density profile ($t=40$) and (b) the time evolution of the density profile of the simulation for $\omega=33$, $k_{\mathrm{dc}}=-0.56$, $k_{\mathrm{ac}}=27.1$, $t_{\mathrm{p}}=8\pi$ and the fraction of atoms in different jets in (c) linear and (d) logarithmic scale. The horizontal line at 0.006 represents the threshold used to construct the phase diagram.
Figure 3: (a) Phase diagram of jets from GPE simulations. Different colors represent regions where different types of jets are above the threshold. (b) Final density profiles in momentum space $|\psi(q)|^{2}$ ($t=8\pi$) for $\omega=33$ [along the dotted line in (a)]. (c) and (e) Experimental phase diagram of jets, measured at $t_{p}=$ 30 ms and 10 ms. The grey lines indicate where the number of atoms falls by 25% and 50% from the initial number of atoms in the condensate. The dashed black line in (c) is a fit to the boundary between the ”no jets” and ”J1” region, proportional to $\omega^{1/2}$. The error bars correspond to the variance of the atom number and the precision of determining the amplitude corresponding to the boundary between the regions. (d) and (f) Averages of 50 absorption images as examples of different points in the phase diagram (marked with black crosses in the diagrams above). From bottom to top the images represent regions with: no jets, only J1, J1 and J2, and J1, J2 and G2 (top two images). The images were taken for evolution time $t_{e}=$ 42.7 ms. $k_{\mathrm{ac}}=Na_{\mathrm{ac}}/a_{r}$, where $N=4000\pm 200$ is the number of atoms in the BEC, before we start modulation. $a_{\mathrm{ac}}$ is calculated from the amplitude of the pulse by a calibration determined using the Feshbach resonance (see Methods).
@@ I.2 Numerical simulation of Gross-Pitaevskii equation
Based on calculated atom fractions, we construct a phase diagram of visibility of different sets of jets, shown in Fig. 3(a). A new region is declared to exist when the fraction for the next jet exceeds 0.006. We choose this value so that it is much larger than the noise in the simulation, but smaller than the atom fraction of the jets at high modulation amplitudes. As the amplitude $k_{\mathrm{ac}}$ increases, additional jets appear, represented by new regions in the phase diagram. As discussed above, the transitions are actually continuous and not discrete, except from the “no jets” to the “J1” region. G3 jets occur in parts of the region marked with ”J1+J2+J3+G2” in Fig. 3(a).
@@ I.3 Experimental phase diagram of jet formation
In contrast to simulation, in the experiments, the number of atoms in jets varies from shot to shot; for example, see Fig. 1(a). This was previously observed in 2D [27], where the distribution of population in angular modes resembles the distribution of thermal radiation. More information on the statistics of the number of atoms in each type of jet is available in the Supplemental Material (Fig. S3). Because of this variance, gathering statistics is necessary to determine the presence of the phase at each point in the phase diagram. We choose the boundary to be at the point where at least one image out of ten shows the corresponding jets. Examples of averages of 50 absorption images of jets from different parts of the phase diagram are shown in Fig. 3(d) and (f) and in the Supplemental Material (Fig. S2). Both Fig. 3(d) and (f) show experimental images for modulation frequency 3 kHz, but (f) is for the shorter modulation time $t_{p}=$ 10 ms and (d) is for $t_{p}=$ 30 ms. Note that in Fig. 3(d) the distance between the jets and the initial condensate varies for different modulation amplitudes. This is because jets are emitted sooner if the modulation amplitude is higher (see Supplemental Material Fig. S6). This effect is not as pronounced Fig. 3(f) due to the shorter modulation time.
The range of frequencies in the experimental and the theoretical phase diagram are similar (the frequencies shown in Fig.  3(a) range from 0 to 7.2 kHz (from 0 to 80 in units of $\omega_{r}=2\pi\cdot 90\mathrm{Hz}$), whereas the experimental phase diagrams in Fig.  3(c) and (e) were measured for frequencies between 1 kHz and 7 kHz). On the other hand, the values of $k_{\mathrm{ac}}$ required for excitation of jets are approximately two times lower in the experiment than in simulation. The difference is most probably because the simulation describes a true 1D system, while in the experiment, the BEC is a 3D object with jets confined into quasi-1D system by a dipole trap beam.
@@
As we increase the amplitude of modulation in the experiment, the total number of atoms starts decreasing and the incoherent background between the jets becomes stronger. This can be seen in Fig. 3(c) and (e), where we plot the thresholds at which the total number of atoms is 25% and 50% smaller than the number of atoms in the initial BEC. The decrease in the number of atoms is likely due to partial collapses of the BEC, caused by the strong attractive interactions the BEC is subjected to during the modulation, leading to three-body losses. The incoherent background is formed by the excited atoms that are not part of the jets. Their velocities have random directions and that appears as a homogeneous background between J0 and J1, and to a lesser extent between J1 and J2. The momentum distribution of the incoherent background remains of the same width for increasing amplitude and does not increase beyond the position of J2 jets as the supercontinuum would. Atom losses and increased fraction of incoherent background, both not accounted for in the simulation, decrease the effective modulation amplitude and prevent the supercontinuum from forming.
To summarize, the experimental phenomena that the 1D GPE simulation does not account for are the atom losses into the incoherent background and three-body losses. For a complete description of the experiment, one would not only need to take into account the three-dimensionality of the experiment, but also the higher order interaction between atoms (e. g. three-particle interaction) and incoherent processes, that are not described by the mean field picture. Despite all of these differences, the discrepancy between the simulation and the experiment only becomes significant at higher modulation amplitudes.
@@ III.2 Numerical simulation
in the dimensionless form [30], where time $t$ is given in units $1/\omega_{r}$, $\omega$ in units $\omega_{r}$, and distances in units $a_{r}=\sqrt{\hbar/m\omega_{r}}$. $k$ is the interaction parameter $k(t)=Na(t)/a_{r}$. We solve the equation on an uniform 1D grid of 65536 points in the interval [-1200:1200]. The initial state of the BEC is determined by an imaginary-time evolution, while the calculation of the real-time dynamics is performed with a time-step of $\tau=$ 0.001. The code is written in Wolfram Language (Mathematica) and it can solve for the ground state through imaginary-time propagation of an initial approximation, and determine the time-dependence in real-time. Most of the computation time is spent in performing the (inverse) Fourier transform, which is performed highly efficiently in Mathematica, thus the performance of the code is comparable to that of compiled codes while it is significantly more flexible. Further details and examples can be found in Ref. [31].